# 2025-04-30

### v1

@@ 2 Measures of Similarity Between Linear Subspaces
The maximal distance value $\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})$ can render is $\sqrt{max(p,q)}$. This corresponds to a scenario when the subspaces $\mathcal{X}$ and $\mathcal{Y}$ are orthogonal to each other, resulting in $\tilde{D}=0$. We normalize the distance measure by its maximal value and introduce the following:
$\hat{\mathcal{D}}_{2}(\mathcal{X},\mathcal{Y})=\frac{\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})}{\sqrt{max(p,q)}}=\sqrt{1-\frac{\tilde{\mathcal{D}}^{2}(\mathcal{X},\mathcal{Y})}{\max(p,q)}},$ (2.40)
@@
$\hat{\mathcal{D}}_{2}(\mathcal{X},\mathcal{Y})=\sqrt{1-\frac{\|X^{\top}Y\|_{F}^{2}}{p}}.$ (2.41)
@@
$\hat{\mathcal{D}}_{2}(\mathcal{X},\mathcal{Y})$ $=\sqrt{1-\frac{\sum_{k=1}^{p}(\cos\theta_{k})^{2}}{p}}=\frac{1}{\sqrt{p}}\sqrt{\sum_{k=1}^{p}\left(1-(\cos\theta_{k})^{2}\right)},$ (2.42)
@@ 3 Subspace-Distance-Enabled Active Learning Framework
$P^{*}=\{\bm{\bar{\mu}}_{1},\bm{\bar{\mu}}_{2},\dots,\bm{\bar{\mu}}_{m_{2}}\}\ \text{with}\ \bm{\bar{\mu}}_{j}\neq\bm{\mu}_{i};\ i=1,\dots,m_{1};\ j=1,\dots,m_{2}.$ (3.4)
@@
$\|\bm{\mu}_{m}-\bm{\mu}_{n}\|_{2}\text{, where }\bm{\mu}_{m}\in P,\ \bm{\mu}_{n}\in P,\ \bm{\mu}_{m}\neq\bm{\mu}_{n}.$ (3.9)
The parameter values with the least Euclidean distance for each sample $\bm{\mu}_{i}\in P$ can be identified, and pairs of nearest neighbors be formed. For instance, when $N_{\mu}=1$, and the values in the set $P$ are ordered in an increasing fashion, each scalar parameter $\mu_{i}\in P$ will only have a maximum of two neighboring parameters $\mu_{i-1}$ and $\mu_{i+1}$. The minimum and maximum $\mu_{i}$ will only have $\mu_{i+1}$ and $\mu_{i-1}$ as the immediate neighbors, respectively. Furthermore, since the size of set $P$ is $m_{1}$, the total number of pairs become $(m_{1}-1)$. For $N_{\mu}>1$, instead of taking the brute-force approach of checking the Euclidean distance between every possible combination of $\bm{\mu}_{i}\in P$, we can resort to a quicker nearest-neighbor lookup using kd-tree [25, 6, 7, 45].
@@
2:  Construct $\mu_{i}$-specific matrix of basis vectors $\Phi(\bm{\mu}_{i})\ \forall\bm{\mu}_{i}\in P$ by following eqs. 3.7 and 3.8.
@@
2:  Construct $\mu_{i}$-specific matrix of basis vectors $\Phi(\bm{\mu}_{i})\ \forall\bm{\mu}_{i}\in P$ by following eqs. 3.7 and 3.8.
@@
The parameter set $P$ is extended by including the selected candidate parameter sample $\bm{\bar{\mu}}^{(iter)}$. At the same time, the candidate set is also updated, $P^{*}=P^{*}\backslash\bm{\bar{\mu}}^{(iter)}$. We query the full-order model solver at the new sample and obtain the snapshots $U(\bm{\bar{\mu}}^{(iter)})$. Next, POD of the new snapshot matrix is carried out, and an appropriate matrix $\Phi(\bm{\bar{\mu}}^{(iter)})$ is constructed by following eq. 3.7, such that the same energy criterion $\eta$, refer eq. 3.8, is maintained for the new sample as well. The neighboring parameter pairs that correspond to the samples in the modified set $P$ are identified, which also results in a change of the total number of pairs $n_{p}$. At the same time, the set of index-pairs is also updated by using eq. 3.10. For $iter>1$, only the distances between any new pairs $(\bm{\mu}_{k},\bm{\mu}_{l})\in I_{pair}$ is evaluated by using either eq. 3.11 or eq. 3.12. The procedures in eqs. 3.13 and 3.14, and that of extending P, are then repeated till a stopping criterion is met.
In Algorithms 3.1 and 3.2, we present two variants to systematically terminate the iterative active learning procedure. One approach is to conclude the selection of new parameters as soon as a user-defined computational budget for the FOM solver queries is met. This quantity is denoted by $max\_query$ in Algorithm 3.1. Another approach is to conclude the selection of new parameters as soon as the user-defined tolerances for the subspace distance and ROM error are satisfied. These tolerances are denoted by $tol_{d}$ and $tol_{e}$ in Algorithm 3.2. The maximal distance between the subspaces is stored as $\mathcal{D}_{max}^{(iter)}$ at steps $5$ and $14$ of Algorithm 3.2. In this variant, the iterations continue as far as $\mathcal{D}_{max}^{(iter)}>tol_{d}$, after which the error induced in the ROM, $\mathcal{E}^{(iter)}$, is evaluated, and the new parameter selection is finally concluded as soon as $\mathcal{E}^{(iter)}\leq tol_{e}$.
@@
In principle, instead of picking new parameter samples by following eq. 3.14, we can simply consider the midpoint of the line-segment connecting $\bm{\mu}_{a}$ to $\bm{\mu}_{b}$ as the new parameter sample, i.e., $\bm{\bar{\mu}}^{(iter)}=\frac{\bm{\mu}_{a}+\bm{\mu}_{b}}{2}$. However, we intentionally avoid this strategy and pick new samples from the candidate parameter set $P^{*}$ to avoid repeated selection of samples in a small sub-region of the parameter domain. Since $P^{*}$ consists some finite number of candidate parameter locations, it implicitly provides a restriction for new sample selection beyond a certain closeness in the parameter domain. As a result, whenever there are no candidate parameter locations left in the region surrounded by $\bm{\mu}_{a}$ and $\bm{\mu}_{b}$, automatically, the new sample can then be selected by using the parameter pair whose subspaces are second farthest. Such a scenario can progressively occur for any parameter pair under consideration, ultimately leading to the exhaustion of all the candidate parameter locations.
@@ 4.1 POD-KSNN ROM
where, $y_{k}(\bm{{x}}_{j})=\widehat{y}_{k}(\bm{{x}}_{j})$ is enforced for all $j=1,\ldots,\ell$ and $k=1,\ldots,q$. Thereby, the network can be trained via the alternating dual-staged iterative training (ADSIT) procedure from [32], either in interpolation mode ($n_{c}=\ell$) or in regression mode ($n_{c}<\ell$). Upon training, we obtain the output layer weights $\{w_{i}^{(k)}\}_{i=1}^{\ell}$ corresponding to $k=1,\ldots,q$, the kernel widths $\{\epsilon_{i}\}_{i=1}^{n_{c}}$, and/or the kernel center locations $\{\mathbf{c}_{i}\}_{i=1}^{n_{c}}$.
@@
$\widehat{\mathbf{y}}_{u}(\bm{\mu}_{i})=\left[\begin{matrix}\mathbf{u}(t_{0},\bm{\mu}_{i})\\ \mathbf{u}(t_{1},\bm{\mu}_{i})\\ \vdots\\ \mathbf{u}(t_{N_{t}},\bm{\mu}_{i})\end{matrix}\right].$ (4.2)
@@
In the online phase of POD-KSNN, $\mu$-KSNN is queried at a new parameter $\bm{\mu^{*}}$ and a SVD of the interpolated snapshots $U^{I}(\bm{\mu}^{*})$ is performed, similar to eq. 3.6,
$U^{I}(\bm{\mu}^{*})=L(\bm{\mu}^{*})\Sigma(\bm{\mu}^{*})K^{\top}(\bm{\mu}^{*}).$ (4.5)
@@
where the truncation level $r_{*}$ is the lowest value such that at least a user-defined energy criterion $\hat{\eta}$ is satisfied, as defined below:
$\hat{\eta}=1-\frac{\sum_{k=1}^{k=r_{*}}(\sigma_{k})^{2}}{\sum_{k=1}^{k=s_{*}}(\sigma_{k})^{2}}.$ (4.7)
@@
Here, $\alpha^{j}\in\mathbb{R}^{r_{*}}$ represents the reduced solution at time $t_{j}$; $r_{*}$ is also known as the size of the ROM. To obtain the reduced solution at any arbitrary time, an interpolant $t$-KSNN is formed in the time domain by following eq. 4.1,
@@ 4.1.1 SDE-ActLearn-POD-KSNN ROM
The subspace-distance-enabled active-learning-driven POD-KSNN (SDE-ActLearn-POD-KSNN) ROM can be constructed as outlined in Algorithms 4.3 and 4.4. The ROM is constructed by iteratively sampling new locations in the parameter domain using either Algorithm 3.1 or Algorithm 3.2 and then constructing the $\mu$-KSNN surrogate. Two variants of the algorithm—one where the ROM is constructed such that a predefined computational budget is met, and other where the ROM is constructed such that a user-defined distance and error tolerances are met—are outlined in Algorithms 4.3 and 4.4, respectively. The online stage of SDE-ActLearn-POD-KSNN ROM stays the same as Algorithm 4.2.
@@
0:  Initial parameter set P, $U(\bm{\mu}_{i})\ \forall\bm{\mu}_{i}\in P$, candidate parameter set $P^{*}$, energy criterion $\eta$, maximum permissible iterations $max\_iter$.
@@ 4.2 POD-NN ROM
The POD-NN method outlined in this subsection is as presented in [59] for parametrized unsteady dynamical systems. The original variant of the method first appeared in [28], which is only applicable to parametrized steady-state systems. To make the method applicable for time-dependent problems, authors in [59] resort to a two-step POD procedure. A temporal compression of the snapshots is done by performing a parameter-specific POD, which is then followed by a compression in the parameter space by applying POD to the collection of parameter-specific POD bases. The final set of dominant POD modes are then used to represent the parametric FOM snapshots in the reduced coordinates. A regression model is constructed for the reduced coordinates using a feedforward neural network, which can be queried at out-of-training time and parameter values. The high-dimensional solution at the new time and parameter value is obtained by using the predicted reduced coordinates and the final set of dominant POD modes.
@@
$U(\bm{\mu}_{i})=[\ \mathbf{u}(t_{0},\bm{\mu}_{i})\ |\ \mathbf{u}(t_{1},\bm{\mu}_{i})\ |\ \dots\ |\ \mathbf{u}(t_{N_{t}},\bm{\mu}_{i})\ ],$ (4.12)
@@
Here, the truncation level $r_{i}$ is the lowest value such that at least a user-defined energy criterion $\hat{\eta}$ is satisfied:
@@
where $s_{i}$ are the total number of nonzero singular values sitting on the main diagonal of $\Sigma(\bm{\mu}_{i})$.
@@
where $\tilde{L}=[\tilde{\mathbf{l}}_{1}|\tilde{\mathbf{l}}_{2}|\dots|\tilde{\mathbf{l}}_{N}]\in\mathbb{R}^{N\times N}$, $\tilde{K}\in\mathbb{R}^{(\sum_{i=1}^{m}r_{i})\times(N_{t}+1)}$ are orthogonal matrices with their columns as the left and right singular vectors respectively. The non-zero singular values in eq. 4.16 are denoted by $\tilde{\sigma}_{k}$, $k=1,\dots,s_{global}$. The matrix $\mathbb{V}$ of reduced basis is formed by selecting the first $r_{global}\leq s_{global}$ dominant POD modes,
@@
$B(\bm{\mu}_{i})=\mathbb{V}^{\top}U(\bm{\mu}_{i})=[\ \beta^{0}(\bm{\mu}_{i})\ |\ \beta^{1}(\bm{\mu}_{i})\ |\ \dots\ |\ \beta^{N_{t}}(\bm{\mu}_{i})\ ],$ (4.19)
where $\beta^{j}\in\mathbb{R}^{r_{global}}$ is the reduced solution at time $t_{j}$. To obtain the reduced solution at any arbitrary time and parameter instance, a feed-forward neural network surrogate is trained using the data $\{\beta^{j}(\bm{\mu}_{i})\}_{j=0}^{N_{t}}$ for all $\bm{\mu}_{i}\in\hat{P}$, and provides an NN-approximation for the reduced solution. The neural network surrogate generates a map $\hat{\beta}$,
$\hat{\beta}:[0,T]\times\mathcal{P}\subset\mathbb{R}^{N_{\mu}+1}$ $\rightarrow\mathbb{R}^{r_{global}},$ (4.20)
$(t,\bm{\mu})$ $\rightarrow\mathbb{V}^{\top}\mathbf{u}(t,\bm{\mu}).$ (4.21)

### v2

@@ 2 Measures of Similarity Between Linear Subspaces
The maximal distance value $\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})$ can render is $\sqrt{\max(p,q)}$. This corresponds to a scenario when the subspaces $\mathcal{X}$ and $\mathcal{Y}$ are orthogonal to each other, resulting in $\tilde{D}=0$. We normalize the distance measure by its maximal value and introduce the following:
$\hat{\mathcal{D}}_{2}(\mathcal{X},\mathcal{Y})=\frac{\mathcal{D}_{2}(\mathcal{X},\mathcal{Y})}{\sqrt{\max(p,q)}}=\sqrt{1-\frac{\tilde{\mathcal{D}}^{2}(\mathcal{X},\mathcal{Y})}{\max(p,q)}},$ (2.14)
@@
$\hat{\mathcal{D}}_{2}(\mathcal{X},\mathcal{Y})=\sqrt{1-\frac{\|X^{\top}Y\|_{F}^{2}}{p}}.$
@@
$\hat{\mathcal{D}}_{2}(\mathcal{X},\mathcal{Y})$ $=\sqrt{1-\frac{\sum_{k=1}^{p}(\cos\theta_{k})^{2}}{p}}=\frac{1}{\sqrt{p}}\sqrt{\sum_{k=1}^{p}\left(1-(\cos\theta_{k})^{2}\right)},$
@@ 3 Subspace-Distance-Enabled Active Learning Framework
$P^{*}=\{\bm{\bar{\mu}}_{1},\bm{\bar{\mu}}_{2},\dots,\bm{\bar{\mu}}_{m_{2}}\}\ \text{with}\ \bm{\bar{\mu}}_{j}\neq\bm{\mu}_{i};\ i=1,\dots,m_{1};\ j=1,\dots,m_{2}.$
@@
$\|\bm{\mu}_{m}-\bm{\mu}_{n}\|_{2}\text{, where }\bm{\mu}_{m}\in P,\ \bm{\mu}_{n}\in P,\ \bm{\mu}_{m}\neq\bm{\mu}_{n}.$
The parameter values with the least Euclidean distance for each sample $\bm{\mu}_{i}\in P$ can be identified, and pairs of nearest neighbors be formed. For instance, when $N_{\mu}=1$, and the values in the set $P$ are ordered in an increasing fashion, each scalar parameter $\mu_{i}\in P$ will only have a maximum of two neighboring parameters $\mu_{i-1}$ and $\mu_{i+1}$. The minimum and maximum $\mu_{i}$ will only have $\mu_{i+1}$ and $\mu_{i-1}$ as the immediate neighbors, respectively. Furthermore, since the size of set $P$ is $m_{1}$, the total number of pairs becomes $(m_{1}-1)$. For $N_{\mu}>1$, instead of taking the brute-force approach of checking the Euclidean distance between every possible combination of $\bm{\mu}_{i}\in P$, we can resort to a quicker nearest-neighbor lookup using kd-tree [25, 6, 7, 45].
@@
2:  Construct $\mu_{i}$-specific matrix of basis vectors $\Phi(\bm{\mu}_{i})\ \forall\bm{\mu}_{i}\in P$ by following eqs. 3.5 and 3.6.
@@
2:  Construct $\mu_{i}$-specific matrix of basis vectors $\Phi(\bm{\mu}_{i})\ \forall\bm{\mu}_{i}\in P$ by following eqs. 3.5 and 3.6.
@@
The parameter set $P$ is extended by including the selected candidate parameter sample $\bm{\bar{\mu}}^{(iter)}$. At the same time, the candidate set is also updated, $P^{*}=P^{*}\backslash\bm{\bar{\mu}}^{(iter)}$. We query the full-order model solver at the new sample and obtain the snapshots $U(\bm{\bar{\mu}}^{(iter)})$. Next, POD of the new snapshot matrix is carried out, and an appropriate matrix $\Phi(\bm{\bar{\mu}}^{(iter)})$ is constructed by following eq. 3.5, such that the same energy criterion $\eta$, refer eq. 3.6, is maintained for the new sample as well. The neighboring parameter pairs that correspond to the samples in the modified set $P$ are identified, which also results in a change of the total number of pairs $n_{p}$. At the same time, the set of index-pairs is also updated by using eq. 3.7. For $iter>1$, only the distances between any new pairs $(\bm{\mu}_{k},\bm{\mu}_{l})\in I_{pair}$ are evaluated by using either eq. 3.8 or eq. 3.9. The procedures in eqs. 3.10 and 3.11, and that of extending P, are then repeated till a stopping criterion is met.
In Algorithms 3.1 and 3.2, we present two variants to systematically terminate the iterative active learning procedure. One approach is to conclude the selection of new parameters as soon as a user-defined computational budget for the FOM solver queries is met. This quantity is denoted by $max\_query$ in Algorithm 3.1. Another approach is to conclude the selection of new parameters as soon as the user-defined tolerances for the subspace distance and ROM error are satisfied. These tolerances are denoted by $tol_{d}$ and $tol_{e}$ in Algorithm 3.2. The maximal distance between the subspaces is stored as $\mathcal{D}_{max}^{(iter)}$ at steps $5$ and $14$ of Algorithm 3.2. In this variant, the iterations continue as long as $\mathcal{D}_{max}^{(iter)}>tol_{d}$, after which the error induced in the ROM, $\mathcal{E}^{(iter)}$, is evaluated, and the new parameter selection is finally concluded as soon as $\mathcal{E}^{(iter)}\leq tol_{e}$.
@@
In principle, instead of picking new parameter samples by following eq. 3.11, we can simply consider the midpoint of the line-segment connecting $\bm{\mu}_{a}$ to $\bm{\mu}_{b}$ as the new parameter sample, i.e., $\bm{\bar{\mu}}^{(iter)}=\frac{\bm{\mu}_{a}+\bm{\mu}_{b}}{2}$. However, we intentionally avoid this strategy and pick new samples from the candidate parameter set $P^{*}$ to avoid repeated selection of samples in a small sub-region of the parameter domain. Since $P^{*}$ contains some finite number of candidate parameter locations, it implicitly provides a restriction for new sample selection beyond a certain closeness in the parameter domain. As a result, whenever there are no candidate parameter locations left in the region surrounded by $\bm{\mu}_{a}$ and $\bm{\mu}_{b}$, automatically, the new sample can then be selected by using the parameter pair whose subspaces are second farthest. Such a scenario can progressively occur for any parameter pair under consideration, ultimately leading to the exhaustion of all the candidate parameter locations.
@@ 4.1 POD-KSNN ROM
where $y_{k}(\bm{{x}}_{j})=\widehat{y}_{k}(\bm{{x}}_{j})$ is enforced for all $j=1,\ldots,\ell$ and $k=1,\ldots,q$. Thereby, the network can be trained via the alternating dual-staged iterative training (ADSIT) procedure from [32], either in interpolation mode ($n_{c}=\ell$) or in regression mode ($n_{c}<\ell$). Upon training, we obtain the output layer weights $\{w_{i}^{(k)}\}_{i=1}^{\ell}$ corresponding to $k=1,\ldots,q$, the kernel widths $\{\epsilon_{i}\}_{i=1}^{n_{c}}$, and/or the kernel center locations $\{\mathbf{c}_{i}\}_{i=1}^{n_{c}}$.
@@
$\widehat{\mathbf{y}}_{u}(\bm{\mu}_{i})=\left[\begin{matrix}\mathbf{u}(t_{0},\bm{\mu}_{i})\\ \mathbf{u}(t_{1},\bm{\mu}_{i})\\ \vdots\\ \mathbf{u}(t_{N_{t}},\bm{\mu}_{i})\end{matrix}\right].$
@@
In the online phase of POD-KSNN, $\mu$-KSNN is queried at a new parameter $\bm{\mu^{*}}$ and the SVD of the interpolated snapshots $U^{I}(\bm{\mu}^{*})$ is performed, similar to eq. 3.4,
$U^{I}(\bm{\mu}^{*})=L(\bm{\mu}^{*})\Sigma(\bm{\mu}^{*})K^{\top}(\bm{\mu}^{*}).$
@@
where the truncation level $r_{*}$ is the lowest value such that the user-defined energy criterion $\hat{\eta}$ is satisfied, as defined below:
$\hat{\eta}=1-\frac{\sum_{k=1}^{k=r_{*}}(\sigma_{k})^{2}}{\sum_{k=1}^{k=s_{*}}(\sigma_{k})^{2}}.$
@@
Here, $\alpha^{j}\in\mathbb{R}^{r_{*}}$ represents the reduced solution at time $t_{j}$. To obtain the reduced solution at any arbitrary time, an interpolant $t$-KSNN is formed in the time domain by following eq. 4.1,
@@ 4.1.1 SDE-ActLearn-POD-KSNN ROM
The subspace-distance-enabled active-learning-driven POD-KSNN (SDE-ActLearn-POD-KSNN) ROM can be constructed as outlined in Algorithms 4.3 and 4.4. The ROM is constructed by iteratively sampling new locations in the parameter domain using either Algorithm 3.1 or Algorithm 3.2 and then constructing the $\mu$-KSNN surrogate. Two variants of the algorithm—one where the ROM is constructed such that a predefined computational budget is met, the other where the ROM is constructed such that a user-defined distance and error tolerances are met—are outlined in Algorithms 4.3 and 4.4, respectively. The online stage of SDE-ActLearn-POD-KSNN ROM stays the same as Algorithm 4.2.
@@
0:  Initial parameter set $P$, $U(\bm{\mu}_{i})\ \forall\bm{\mu}_{i}\in P$, candidate parameter set $P^{*}$, energy criterion $\eta$, maximum number of permissible iterations $max\_iter$.
@@ 4.2 POD-NN ROM
The POD-NN method outlined in this subsection is as presented in [59] for parametrized dynamical systems. The original variant of the method first appeared in [28], which is only applicable to parametrized steady-state systems. To make the method applicable for time-dependent problems, authors in [59] resort to a two-step POD procedure. A temporal compression of the snapshots is done by performing a parameter-specific POD, which is then followed by a compression in the parameter space by applying POD to the collection of parameter-specific POD bases. The final set of dominant POD modes is then used to represent the parametric FOM snapshots in the reduced coordinates. A regression model is constructed for the reduced coordinates using a feedforward neural network, which can be queried at out-of-training time and parameter values. The high-dimensional solution at the new time and parameter value is obtained by using the predicted reduced coordinates and the final set of dominant POD modes.
@@
$U(\bm{\mu}_{i})=[\ \mathbf{u}(t_{0},\bm{\mu}_{i})\ |\ \mathbf{u}(t_{1},\bm{\mu}_{i})\ |\ \dots\ |\ \mathbf{u}(t_{N_{t}},\bm{\mu}_{i})\ ],$
@@
Here, the truncation level $r_{i}$ is the lowest value such that the user-defined energy criterion $\hat{\eta}$ is satisfied:
@@
where $s_{i}$ is the total number of nonzero singular values sitting on the main diagonal of $\Sigma(\bm{\mu}_{i})$.
@@
where $\tilde{L}=[\tilde{\mathbf{l}}_{1}|\tilde{\mathbf{l}}_{2}|\dots|\tilde{\mathbf{l}}_{N}]\in\mathbb{R}^{N\times N}$, $\tilde{K}\in\mathbb{R}^{(\sum_{i=1}^{m}r_{i})\times(N_{t}+1)}$ are orthogonal matrices with their columns as the left and right singular vectors, respectively. The non-zero singular values in eq. 4.11 are denoted by $\tilde{\sigma}_{k}$, $k=1,\dots,s_{global}$. The matrix $\mathbb{V}$ representing the reduced basis is formed by selecting the first $r_{global}\leq s_{global}$ dominant POD modes,
@@
$B(\bm{\mu}_{i})=\mathbb{V}^{\top}U(\bm{\mu}_{i})=[\ \beta^{0}(\bm{\mu}_{i})\ |\ \beta^{1}(\bm{\mu}_{i})\ |\ \dots\ |\ \beta^{N_{t}}(\bm{\mu}_{i})\ ],$
where $\beta^{j}\in\mathbb{R}^{r_{global}}$ is the reduced solution at time $t_{j}$. To obtain the reduced solution at any arbitrary time and parameter instance, a deep feed-forward neural network surrogate is trained using the data $\{\beta^{j}(\bm{\mu}_{i})\}_{j=0}^{N_{t}}$ for all $\bm{\mu}_{i}\in\hat{P}$, providing an NN-approximation for the reduced solution. The deep neural network surrogate generates a map $\hat{\beta}$,
$\hat{\beta}:[0,T]\times\mathcal{P}\subset\mathbb{R}^{N_{\mu}+1}$ $\rightarrow\mathbb{R}^{r_{global}},$
$(t,\bm{\mu})$ $\rightarrow\mathbb{V}^{\top}\mathbf{u}(t,\bm{\mu}).$